OpenStudy (anonymous):

If w=⟨–7,–3⟩ what is w · w? 40 40√ 58 58√

OpenStudy (raden):

if w = (a1, b1) then w. w = a1^2 + b1^2 now plug in a1=-7 and b1 = -3

OpenStudy (anonymous):

i ogt -58?

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

negative is squared always be postive (-7)^2 + (-3)^2 = 49 + 9 = 58